OpenStudy (anonymous):

write an equation of the line that intersects the y-axis at (0,1), and is perpendicular to the line whos equation is x+2y=6

OpenStudy (saifoo.khan):

x+ 2y = 6 FInd it's slope. THen find the slope of the line perpendicular to it using, \[m_1 \times m_2 = -1\] FInally use the point given and the slope to find the equation

OpenStudy (anonymous):

slope would be 2 for the first one right

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

-1/2*m=-1 m=2

OpenStudy (saifoo.khan):

Nope, solve for y. x+ 2y = 6 2y = -x + 6 y = -1/2 x + 3 So slope = -1/2 Now find the slope of the line perp to it using the formula given above.

OpenStudy (anonymous):

y=2x-1?

OpenStudy (saifoo.khan):

Perfect.

OpenStudy (anonymous):

thanks bro.

OpenStudy (saifoo.khan):

Welcome.